Subject: APO 3.5m users committee minutes, 2/28/11

From: Suzanne Hawley

Submitted: Mon, 7 Mar 2011 15:51:38 -0800 (PST)

Message number: 1219 (previous: 1218, next: 1220 up: Index)

 		APO 3.5-m Users Committee Phonecon, 2/28/11

Attending:  Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback, Remy 
Indebetouw, Sean Moran, Jon Holtzman, Guy Stringfellow (for John Bally), Dmitry 
Bizyaev, Bruce Gillespie

Absent: Al Harper


User feedback, comments from institutional representatives:

o Chicago (Al Harper) - absent.

o New Mexico State (Jon Holtzman) - Jon had nothing to report.

o Washington (Scott Anderson) - Scott had nothing to report.

o Colorado (Guy Stringfellow) - Guy mentioned that the recent changes to the 
image headers are a good thing, and it might also be useful to look at header 
information in general to see if we're capturing everything that is needed, 
e.g., eyelid status. There was some discussion, but Mark pointed out that the 
eyelid status is not accessible to TUI by design; you can use a line command to 
see if an eyelid has been commanded to move, but its actual state and history 
are not archived.

o Princeton (Michael Strauss) - Michael said everything is fine.

o Johns Hopkins (Sean Moran) - Sean said that one JHU user was having 
difficulty finding current information on exposure calculations for the 
echelle.  Jon Holtzman mentioned that we used to have an exposure calculator in 
the old documentation, but it had problems and was excised in the current 
version.  Jon said he would think about furnishing an update.

o Virginia (Remy Indebetouw) - Remy said that there are no user problems. 
There is one user that uses Windows TUI.  Remy said he's still working on 
examples of TripleSpec data that demonstrate the World Time System asynchrony 
problem in the TripleSpec headers.


Discussion of telescope/instruments report - Klaene:

Mark said there were no major highlights to report this month.  He mentioned 
the warmup of the DIS blue camera due to a Cryotiger glitch related to the 
recent cold snap.  New software and an upgraded network switch were installed 
to help reduce or eliminate the spontaneous Agile aborts.  The Tspec 
slit-viewer mask was updated.  A cable icing problem caused a slight warm-up of 
the echelle, and we're monitoring its vacuum level and CCD temperature. 
NIC-FPS is back at APO, and is cold and undergoing testing (see separate 
report, below).  The Integral Field Spectrograph (IFS) upgrade to the GFP was 
started, and the modified instrument saw first light briefly and got its 
instrument block updated -- the Goddard team will be back in March to continue 
working on the GFP upgrade, and more will be reported at our next meeting.

Guy asked whether updates to slit masks in general should be reflected in the 
user information.  Bill said that the artifacts in the image remain unchanged,
and that updates usually address the vignetting correction.

Mark's detailed report for the last month follows:


          3.5-m Telescope, Instruments, and CIF Projects Highlights
                         1/27/10 through 2/23/11

0) Overview

Weather has been generally good barring one large storm system that deposited 
6" of snow and brought severely cold temperatures not only to this area but 
most of the continental US.  So far, only minor problems occurred as a result 
of these severe temperatures. However, we believe there is a water leak 
somewhere in our underground plumbing system. The forest remains dry and 
extreme forest fire danger is expected later this spring.

1) Telescope

Virtually no telescope problems were reported during this period.

A new version of TUI (1.8.6) is available from the usual location:

 	(Mac and unix versions now, Windows to follow)

See the version history for details:


2) Instruments

DIS performed well this period.  The new slit-viewer shutter has been working 
well in the cold weather.  Spectrograph and slit-viewer were refocused.  DIS 
blue camera did warm as a result of the compressor failing to recover from a 
power glitch during a cold day.  The camera re-cooled quickly but had to be 
vacuum pumped shortly thereafter. We are continuing to work on DIS focus issues 
with the slit viewer.

SPICam operated with no new issues.

Agile spontaneous aborts are being worked on.  New kernel and driver were 
installed. Agile was working well with the new system in continuous runs of 
exposures of more than 2 seconds.  Less than 2 second exposures still causes 
some failures and is being worked with an upgrade to the network, among other 
things.  We are proceeding with plans to move the dark slide to the filter 
wheel which when combined with the port eyelid should provide very good darks.

TripleSpec is not reporting slit position in the headers.  This is on the list 
of things to fix.  An issue arose with a dark spot on the slit-viewer processed 
images that was corrected with an updated mask.

The Echelle had an inadvertent warmup of the shield due to a failure of the 
splash sensor.  Both LN2 splash sensors have been retrofitted with a more 
advanced and robust design.  The warmup did not cause any discernible CCD 
contamination however the CCD is running slightly warmer than normal and we 
expect to have to vacuum pump the instrument in the summer, or possibly sooner.

NIC-FPS is being returned to the observatory on Feb 24.  We will be obtaining 
new noise characteristics and will update on further status at the telecon.

GFP upgrade has started.  After 3 weeks at the observatory the new detector and 
IFS optics were installed.  We were able to get an instrument block and test 
basic imaging function but more work is scheduled to work on the software, get 
the IFS working, and hopefully the photon-counting mode operational.  GFP will 
not be available for normal operations during Q2 2011.


NIC-FPS update - Klaene:

NIC-FPS is back at APO, and was pumped and cooled for tests following 
installation of its new fan-out board.  The detector gain was measured at 2.94, 
nearly unchanged. On single reads, the noise is down to 30 electrons, nearly 
three times better than before the new fan-out board was installed.  In NSF 
mode, the noise is about 12 electrons. The linearity is good to 3% up to about 
54k DSN, and 5% beyond.  The NIC-FPS user documentation will be updated with 
these measurements.  Suzanne gave her appreciation to the collaborators at UVa 
and Stephane Beland (CU) for their work on getting us to this

The problem with NIC-FPS data corruption when the instrument is rotated was not 
reproducible when the instrument was in Colorado, and is still with us. 
Stephane and APO staff are trying to isolate and identify the source of this 
intermittent problem, which occurs mostly under slow tracking speeds when the 
cooled instrument is upside down, plus or minus about 60 degrees rotation.  We 
hope to find and fix this problem shortly, or develop a software or procedural 
workaround.  Suzanne added that NIC-FPS will be returned to scheduled service 
in March, pending the results of the ongoing tests.


DIS logging script - Ketzeback:

Bill reported that Adam Kowalski (UW) wrote an IDL logging script that runs on 
3.5m observing computer, newton.  It queries DIS users for their program ID and 
UT date of observations, and then the program looks up the header information 
for DIS science and slit-viewer exposures for that night, and writes a log of 
science images, slit-view images, date of mid-time for science exposures, and 
the deltas between the mid-times of science exposures and relevant slit-view 
images.  Also included in the log is the range of slit-view images taken during 
a science exposure, or the single corresponding slit-view image, or reports 
that there is no slit-view image. The software is nearly ready to be released, 
and we plan to make it available to all DIS users; the DIS user documentation 
will be updated, and notification will be made to the schedulers, users 
committee members and apo35-general.  Suzanne added that we can't easily put 
this information in the exposure headers because of cross-platform
and timing issues.  Remy asked if we could do this for TripleSpec exposures, 
too. Suzanne said it should be possible, if someone wants to spend a little 
time on it
(volunteers welcome).  Remy said he would talk with Matt Nelson and John Wilson 
at UVa about the information currently written into the Tspec headers, and also 
about the synchrony issue between the time recorded for the science and slit 


Future of ARCSAT - Hawley, Bizyaev:

Suzanne announced that Dmitry Bizyaev is the newly appointed Director of the 
0.5-m ARCSAT telescope (previously known at the Photometric Telescope - PT - 
SDSS-I/II), succeeding Stephanie Snedden who was ARCSAT Director the past two 
years. Suzanne was recently approached by a small funding entity and invited to 
apply for a hardware upgrade grant which could conceivably be used for ARCSAT.
ARCSAT is currently being used by some of the APO staff for research projects, 
and also by some other ARC users.  It receives modest support from the 
discretionary line in Suzanne's operations budget, and from volunteer labor 
from site staff. The telescope is not easy to operate remotely, and is still 
running on its original operating system.  It does have an upgraded imaging 
camera useful for short exposures
with good blue-sensitivity that was provided by UW, and is being used to obtain 
simultaneous photometric data with spectroscopy taken with the 3.5-m telescope. 
We have several ideas for upgrading the telescope, including enabling full 
remote operations, an imaging camera for wide-field use (to replace the old PT 
camera), continuing maintenance, and even full robotic operation.  Before we 
proceed with a proposal, we would like to see what the broader ARC community's 
interests would be for scientific utilization of ARCSAT.

Dmitry gave some additional background.  Two groups, one from APO and one from 
UW, are currently splitting the observing time on ARCSAT.  It can be operated 
remotely, but needs on-site intervention for opening and closing, and for 
troubleshooting. We probably need of order $20-30K to make the telescope fully 
remotely operable; DFM (the original manufacturer of the telescope) may be able 
to do this upgrade.  The availability of engineering staff support at the site 
is an issue, because SDSS is no longer using or supporting the telescope.  The 
old SDSS (PT) camera is still working but is difficult to repair or maintain. 
The PT camera has a 40 arcmin FOV, with a 6-slot wheel for 3-inch filters 
(changeable during the day), and long
readout time; the UW camera has a 10 arcmin FOV, a filter wheel with seven 
~2-inch filters, and short readout time.

Jon pointed out that the NMSU 1-m telescope is operated robotically and can 
also be used for obtaining photometric data at APO.  Potential users can 
contact him.

Suzanne proposed the following action item.

 	===> ACTION:  User committee members to poll their users to determine 
would be interested in using ARCSAT, willing to help with an upgrade proposal 
and its implementation, and for what scientific objectives?



[open from previous meeting]:

 	===> ACTION (now closed):  Users Committee members should poll their 
users to find out how many are using TUI with Windows machines and report at 
the next Users Committee meeting.

Apparently Colorado is the only institution that still has a significant number
of Windows users.  Guy said that about a quarter of the CU users use TUI on 
Windows machines.  Remy, Michael, Scott, and Sean all said that their users 
would be OK with cutting support to Windows-based TUI.  Mark said that one of 
the Observing Specialists, Joe Huehnerhoff, was just able to build the current 
version of TUI for Windows and it is
available at the usual location.  For now, TUI will remain available in Windows
as long as Joe is able to support it.

[new actions from this meeting]:

 	===> ACTION:  User committee members to poll their users to determine 
would be interested in using ARCSAT, willing to help with an upgrade proposal 
and its implementation, and for what scientific objectives?


Next meeting:  The next Users Committee phonecon will be on Monday 4 April at 
8:30 AM Pacific Time.  The agenda and other materials will be sent to the 
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1219 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to